OpenStudy (anonymous):

Solve: x(squared) + 4x+8=0

OpenStudy (mimi_x3):

\(x^2+4x+8=0\) \[x=\frac{-(b)\pm\sqrt{(b)^{2}-4(a)(c)}}{2(a)} \]

OpenStudy (anonymous):

First factor the equation, if possible.

OpenStudy (anonymous):

salam22 its a quadratic equation so you want me to make a grafh of it or only zeros?

OpenStudy (anonymous):

If not then you the quadratic equation given by @Mimi_x3

OpenStudy (anonymous):

And in this case b^2-4ac<0 So you will have 2 complex zero's.

OpenStudy (mimi_x3):

Yeah, looks like a complex.

OpenStudy (anonymous):

So just plug in the values a=1, b=4 and c=8. and remember that you should have 2 complex zero's

OpenStudy (anonymous):

The answer choices I have are: A) 2+2i, 2 - 2i B) -2 + 4i, -2 - 4i C) -2+2i, -2 - 2i D) 2+4i, 2-4i Can someone tell me how to do the questions as well as what the answer is?

OpenStudy (anonymous):

Sorry. Actually, I have the answers already. I just don't know how to do the problem.

OpenStudy (anonymous):

That's okay :). Don't be afraid to let us know.

OpenStudy (anonymous):

So,\[x=\frac{-4 \pm \sqrt{(-4)^2-4(1)(8)}}{2*1}\]

OpenStudy (anonymous):

Can you clean that up @salam22

OpenStudy (anonymous):

getting my graphing calculator so I know how to do it. 1 min. please.

OpenStudy (anonymous):

\[x=\frac{-4 \pm \sqrt{16-32}}{2}\] \[x=\frac{-4 \pm \sqrt{16-32}}{2}\] \[x=\frac{-4 \pm \sqrt{-16}}{2}\] \[x=\frac{-4 \pm \sqrt{16}\sqrt{-1}}{2}\] \[\sqrt{-1}=i\] So, \[x=\frac{-4 \pm 4*i}{2}\] \[x=\frac{-4}{2} \pm \frac{4}{2}i\]

OpenStudy (anonymous):

What course is this for? Are you allowed to use a graphing calculator?

OpenStudy (anonymous):

And then am sure you can simplify your solution? You are familiar with complex number right?

OpenStudy (anonymous):

about to log off and log back in using another computer, so my mom can continue using this 1. I have a study guide which my teacher provided the answers, but we're expected to find out how to do the problems. I have a total of 85 questions, and I only need help with 18 of them so I can prepare for my final on Thursday. I'm trying to get the 18 done 2night so I can bring it to my teacher 2morrow. So any help is VERY appreciated. Wish I found this site earlier, cuz i was just about to go to bed, now I have to stay up :-(

OpenStudy (anonymous):

So the answer would be c

OpenStudy (anonymous):

Algebra 2 and yes, calculators are allowed now and during the test

OpenStudy (anonymous):

Oh, I understand... well try using WolframAlpha to see it that also helps because sometimes people take long to reply.

OpenStudy (anonymous):

Well, All the best! Did you understand what I did?

OpenStudy (anonymous):

THANKS SO MUCH! U R THE BOMB!

OpenStudy (anonymous):

Yes, I understand exactly what you did.

OpenStudy (anonymous):

You're surely welcome and...Aww, thanks!

OpenStudy (anonymous):

and btw, I'm not familiar with complex numbers

OpenStudy (anonymous):

Im writing down the answer u provided so I can practice on my own. Will post the next question shortly.

OpenStudy (anonymous):

Oh... yikes... Well, remember how they once said you can't take the square root of a negative number? When raised to an even power. Well you actually can and mathematicians decided to call this set of numbers complex numbers \[i=\sqrt{-1}, i^2=-1, i^3=-i, i^4=1\] and they pretty much repeat the pattern after i^4 started back at i

OpenStudy (anonymous):

And unfortunately I have to go to bed now... but if you post it quickly I can help you with it if I am able to :).

OpenStudy (anonymous):

ok Thank you. solve by completeing the square 8x^2 - 80x = 32. can you help me work this out

OpenStudy (anonymous):

To complete the square you must have the coefficient of the x=1 to 1 so first divide by 8

OpenStudy (anonymous):

I accidently closed this question and started a new one since its a new example. Sorry. Are u still able to reply to this post?

OpenStudy (anonymous):

Okay I will look for it